OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5. y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (anonymous):

y-y1 = m(x-x1) where m is the gradient and (x1,y1) is your point so y+1 = 5(x+2)

OpenStudy (anonymous):

just use the formula y - y1 = m ( x - x1) and put the values y1 = -1,x1 = -2 and m = 5